Given cos(theta)= 1/4 with theta in Quadrant 1, find cos(theta/2) , sin(theta/2) , sin2(theta) , cos2(theta):
Answer by Theo(3464)

cos(t) = 1/4
this makes t = 75.52248781 degrees.
cos(75.52248781) = 1/4
that establishes your baseline angle.
cos(t/2) is equal to cos(75.52248781/2) = .790569415
sin(t/2) is equal to sin(75.52248781/2) = .612372436
sin^2(t) = (sin(t))^2 = (.968245837)^2 = .9375
cos^2(t) = (cos(t))^2 = (25)^2 = .0625
you need to use your calculator to get these values.
find cos^-1(1/4) first and then work from there.